The derivative of (4x)^3 * (2x)^6 is


I know that I have to use the chain rule, but I get lost after that.

Answer by drk(1908)

This is of the form X^a*X^b. we can expand these and rewrite as simply X. This will eliminate chain rule.
(4x)^3 = 64x^3
(2x)^6 = 64x^6
---
(4x)^3 * (2x)^6 = (64^2)(x^9)
y = (64^2)(x^9)
y' = 9*64^2*x^8
or
36864x^8